### v2

# Core-Set Selection for Data-efficient Land Cover Segmentation

Abstract
The increasing accessibility of remotely sensed data and the potential of such data to inform large-scale decision-making has driven the development of deep learning models for many Earth Observation tasks. Traditionally, such models must be trained on large datasets. However, the common assumption that broadly larger datasets lead to better outcomes tends to overlook the complexities of the data distribution, the potential for introducing biases and noise, and the computational resources required for processing and storing vast datasets. Therefore, effective solutions should consider both the quantity and quality of data. In this paper, we introduce six basic core-set selection methods for selecting important subsets of examples from remote sensing image segmentation datasets that rely on imagery only, labels only, and a combination of each. We benchmark these approaches against two traditional baselines on three widely used land-cover classification datasets - DFC2022, Vaihingen, and Potsdam - thus establishing a general and comprehensive baseline for future works. In each of the datasets, we demonstrate that the proposed methods outperform the baselines across various settings, with some approaches even selecting core sets that surpass training on all available data. For example, with the DFC2022 dataset, we find a subset of size 10% that, when used in training, results in a slightly better performing model compared to using the entire dataset. This result shows the importance and potential of data-centric learning for the remote sensing domain. The code is available at https://github.com/keillernogueira/data-centric-rs-classification/.
Index Terms:
Core-set selection, Data-centric machine learning, Land-cover classification, Semantic Segmentation
I Introduction
The rapid advancements in satellite technologies have significantly enhanced the accessibility to Earth observation data, opening new opportunities for a better understanding of the Earth’s surface [1]. Towards this, several deep learning methods have been trained and exploited using increasingly larger labeled data sets [2], often emphasizing label quantity over quality. However, this indiscriminate increase in data volume can lead to diminishing returns or even detrimental effects on model performance, given that the creation of these larger datasets requires significant human effort or integrating weak labels that, in turn, can lead to the introduction of noise, bias, and inaccurate annotations. In general, the current assumption that more data inherently leads to better outcomes may overlook the complexities of data distribution [3], the potential for introducing biases and noise, spurious correlations, the energy consumption [4], and the computational resources required for processing, labeling, and storing vast datasets. Therefore, effective solutions should consider not only the quantity but also the quality of data.
To this end, core-set selection is a technique that focuses on finding and using the most valuable examples for training deep learning models, thus preserving the essential characteristics and insights of the entire dataset while maintaining or even enhancing overall model performance. Such a paradigm can assist in several aspects such as improved computational efficiency, cost-effective data handling, enhanced model performance, effective use of labeled data, or efficient labeling of unlabeled data [5, 6, 7, 8].
Given its importance, several papers have addressed core-set selection, usually employing methods such as clustering algorithms and gradient approximation [9, 10, 11]. Some works perform the core-set selection before training the final machine-learning model [9, 10] whereas others perform a new selection at each training epoch [12, 11, 13]. The former is agnostic regarding the model that is finally applied to the established training data but the selection process is performed only once. On the other hand, the latter is more optimized for the specific machine learning model, but it requires more computational resources since core-set selection is performed multiple times during training. Since we aim to establish a general baseline for core-set selection methods for remote sensing data, we focus on approaches that are independent of the downstream model (i.e., model agnostic), even because these can be easily adapted and used in different scenarios, applications, and tasks, including with modern architectures, such as vision-language and foundation models. Finally, although some of these works perform core-set selection for image classification, to the best of our knowledge, there have been no initiatives exploiting this paradigm for remote sensing image segmentation, which is the focus of our work.
Towards establishing this general and comprehensive baseline, this paper introduces and benchmarks six basic core-set selection approaches for remote sensing image segmentation based on several distinct premises, that rely on imagery only, labels only, and a combination of each. Specifically, given the training instances (i.e., images and their corresponding segmentation masks), the proposed approaches rank the examples from most to least valuable for training, based on the representativeness of each example, according to specific criteria. Instead of training the deep learning-based models using all available data, we leverage the aforementioned ranking to select the most representative examples (core-set) and train the models accordingly. By doing so, we can not only reduce the training time but also improve the overall effectiveness by filtering out non-representative and/or noisy examples. The main contributions of this paper are the following:
  • Introduction of six core-set selection approaches for remote sensing image segmentation, each based on different premises; and
  • A full set of experiments comparing these approaches against two common baselines on three widely used datasets, thus establishing a benchmark for future research in core-set selection.
Overall, this work, an outcome of the Data-Centric Land Cover Classification Challenge of the Workshop on Machine Vision for Earth Observation and Environment Monitoring (MVEO) 2023, fills a critical gap in the literature and demonstrates the potential of core-set selection in advancing remote sensing image segmentation as well as data curation and labeling.
Figure 1: General overview of core-set selection. An input data set is first processed by a core-set selection algorithm that, based on some criteria, prioritizes certain examples over others (represented by the size of the blue circles). Based on this, it is possible to select the core-set data depending on the amount of data one would like to retain (illustrated by red, orange, and green dotted circles). Finally, the selected core-set is used to train a machine learning model, thus reducing the training time while maintaining, or even improving, task performance.
II Related work
One of the major reasons to work with a core-set instead of the full data set is an improvement in computational efficiency. Reducing the size of the dataset allows for quicker processing and experimentation. This makes it possible to use complex machine-learning models without immense computational costs. For this, oftentimes data-only techniques such as naïve random sampling are applied before model training, i.e., these methods do not need access to labels and are independent of the learning objective and application. An example is the identification of a subset that approximates the loss function of the whole dataset [9, 11, 13]. Furthermore, besides computational efficiency, smaller datasets reduce storage and maintenance costs, which is crucial when managing vast amounts of data from Earth observation systems.
Another reason is to improve model performance by enhancing the learning process and reducing overfitting through filtering out noisy or incorrect data points, thus creating cleaner datasets [14]. In remote sensing, predominantly prior knowledge, label information, or an existing model is used to identify a clean core-set [15, 16]. Santos et al. [10], for example, use clustering for satellite time series to identify instances that are mislabeled or have low accuracy, with the goal of removing them from the training set to avoid a decrease in model performance. Moreover, many methods have been developed for data with known sources of uncertainty, such as clouds [17, 18, 19]. Furthermore, this reason is directly related to the enhancement of the accuracy and robustness of machine learning models by removing low-quality or redundant examples. Such models can perform as well as, or even better than, those trained on the full dataset (e.g., [20, 21, 22]).
Another reason for core-set selection is to support clear and non-misleading explanations of a model and the data. Generally, with the field of explainable machine learning, new methods are introduced to calculate importance scores, sensitivities, or contributions of features and interpret them as relevance [23]. However, redundancies and correlations distort the derived insights, therefore they should be removed before interpreting and explaining the results. With the goal to analyze geospatial air quality estimations and the relevance of specific measurement locations, Stadtler et al. [24] demonstrated that removing redundant examples only slightly decreases test accuracy, as these are not relevant for training.
In general, the principles of core-set selection are closely related to areas like active learning [12, 25, 26] and self-training [27]. For unlabeled datasets, core-set selection can guide efficient labeling by identifying the most representative examples. This optimizes resource allocation for manual annotation - a common objective in active learning scenarios [12, 25, 26]. In case the data is already labeled, core-set selection can help prioritize the most informative examples. This maximizes the use of labeled data and may reduce the need for further labeling efforts.
Overall, although most of the aforementioned works perform core-set selection for image classification, to the best of our knowledge, our work is the first research to design and benchmark core-set selection techniques specifically for remote sensing image segmentation.
III Core-set selection methods
Given a set of satellite images $\mathbf{X}$ and their corresponding label masks $\mathbf{M}$ (also known as segmentation maps), we introduce six core-set selection methods that assign an importance score, $s_{i}$, to the $i$-th instance $(X_{i},M_{i})$. These scores range from 0 (least valuable for training) to 1 (most valuable for training). The aim of these methods is to rank the examples based on their informativeness, allowing us to select a subset of the data — a core-set — that can achieve good model performance with reduced training time and data size.
The proposed methods are categorized into: label-based methods, which only rely on the label masks $\mathbf{M}$, image-based methods, which use the information only from the input images $\mathbf{X}$, and combined methods that integrate both sources.
[…]
III-A Label Complexity (LC)
LC is a label-based method that scores an instance based on the complexity of its label mask $M_{i}$. The underlying assumption is that examples with high-complexity label masks are more informative for training segmentation models. It is important to emphasize that this approach does not explicitly guarantee representativeness but instead prioritizes complex label masks to potentially generate more informative training signals, which may lead to improved model performance.
The complexity of a label mask is quantified using the entropy of the class distribution in the label mask – high entropy class distributions will include more and mixed classes, while low or zero entropy class distributions will be dominated by a single class. Precisely, for each instance $i$, we compute the score $s_{i}^{\text{LC}}$ based on the entropy $H(M_{i})$ as:
$s_{i}^{\text{LC}}=H(M_{i})=-\sum_{c=1}^{C}p_{i,c}\log_{C}(p_{i,c}),$ (1)
where $C$ is the number of classes and $p_{i,c}$ is the proportion of pixels belonging to class $c$ in $M_{i}$. Classes that are labeled as “unknown” or “ignored” (in DFC2022) are excluded from this computation.
Higher scores correspond to examples that have a more uniform distribution of class labels with potentially more informative label masks, while low scores correspond to examples that are dominated by a single class.
III-B Feature Space Diversity (FD)
The FD method is an image-based approach that aims to select a diverse core-set of examples from $\mathbf{X}$ based on each example’s embedding from a pre-trained deep learning model.
First, we embed each image, $X_{i}$, using a ResNet-18 model [28] pre-trained on ImageNet [29]. We use the final feature representation layer (i.e. after spatial pooling) from the model, which results in a feature vector, $F_{i}\in\mathbb{R}^{512}$, that encodes higher-level semantic information per image.
Next, we group the feature embeddings into $K$ clusters using the K-Means algorithm. We search for a value of $K$ that minimizes the average Vendi score [30] across clusters. The Vendi score is a measure of diversity over a set of vectors – by minimizing the average within-cluster diversity we ensure that an instance from that cluster is representative of the others. Starting with $K=2$, we cluster all image embeddings with K-Means, measure the average per-cluster Vendi score, then increment $K$, and repeat until the change in the average Vendi score falls below a threshold of $\delta$ (we choose $\delta=0.5\%$) for three iterations.
Given a clustering of the examples, we sequentially choose one instance randomly from within each cluster in a round-robin fashion (the first cluster is randomly selected from the set of $K$ clusters), resulting in ordered $K$-sized segments of cross-cluster examples. The first example has the highest importance score $s_{i}^{\text{FD}}=1$, while the last selected example receives the lowest score $s_{i}^{\text{FD}}=0$.
It is important to highlight here that this random selection step does not inherently produce a fixed ordering of scores. However, this does not impact the primary objective of this method, which is to ensure diversity among the selected examples.
III-C Complexity/Diversity Hybrid (LC/FD)
The LC/FD method is a hybrid method that combines the most important examples from the LC and FD methods, following an assumption that diversity is important when working with small datasets, while label complexity becomes increasingly important for medium to large datasets.
Specifically, the LC/FD method is a hybrid approach that uses the ranked lists of examples from LC and FD, denoted as $\mathcal{R}_{\text{LC}}$ and $\mathcal{R}_{\text{FD}}$, respectively. A cutoff point $m$ is defined, and the hybrid ranking is constructed by taking the top $m$ examples from $\mathcal{R}_{\text{FD}}$, followed by all examples from $\mathcal{R}_{\text{LC}}$ that are not already included. This ensures that the selected core-set includes a mix of label-complex examples and feature-diverse instances.
III-D Feature Activation (FA)
The FA method is an image-based approach that uses statistics from image embeddings created by a pre-trained neural network to rank examples.
First, a ResNet-18 [28], pre-trained on the ImageNet dataset, is used to extract the image embeddings (after the final spatial pooling layer), resulting in a feature vector, $F_{i}\in\mathbb{R}^{512}$, that encodes higher-level semantic information per image.
By construction, all values in $F_{i}$ are non-negative (via application of a ReLU function within the network). We assume that examples with high activation magnitudes (large mean) and significant variations across different dimensions (high standard deviation) in feature space are likely to carry more relevant information. Therefore, we compute the score $s_{i}^{\text{FA}}$ with a combination of the mean $\mu_{i}$ and standard deviation $\sigma_{i}$ (both scaled to (01](0-1]( 0 - 1 ]) of the embedding vector $F_{i}$ as:
$s^{\text{FA}}_{i}=1-\left[\frac{\gamma_{i}-\min\limits_{F_{j}\in\mathbf{F}}[\gamma_{j}]}{\max\limits_{F_{j}\in\mathbf{F}}[\gamma_{j}]-\min\limits_{F_{j}\in\mathbf{F}}[\gamma_{j}]}\right],$ (2)
where $\gamma_{i}=-(1-\mu_{i})\cdot log(\sigma_{i})$.
According to the aforementioned assumption, examples with low scores are likely to have lower diversity, contain more noise, etc, and are therefore likely to be less important for training.
III-E Class Balance (CB)
Similar to the LC, the CB method is a label-based technique that aims to select a subset of examples with a uniform class distribution by using a time-efficient strategy that preprocesses and computes the class distribution of each image for subset selection.
The method consists of $N$ steps, where $N$ refers to the number of examples in the dataset. In each step, the most suitable instance is selected from the dataset to ensure that the overall class distribution of the selected examples approaches a uniform distribution. Specifically, the most suitable instance is the instance that maximizes the entropy of the class distribution of the union of the current core-set with the selected example.
The order in which the examples are selected determines their importance score: the first instance is ranked as the most important and the last example is ranked as least important. Formally, let $r_{i}$ be the rank of instance $i$, then:
$s^{CB}_{i}=1-\frac{r_{i}}{N}.$ (3)
III-F Feature Activation/Class Balance Hybrid (FA/CB)
The FA/CB method is a hybrid method that uses a weighted ensemble of importance scores calculated by the previous two methods, that is:
$s^{\text{FA/CB}}_{i}=\lambda\cdot s^{\text{FA}}_{i}+(1-\lambda)\cdot s^{\text{CB}}_{i},$ (4)
where $\lambda$ is a trade-off weight between the two scores.
IV Experimental Setup
IV-A Datasets
We test our approaches using three high-resolution datasets for semantic segmentation with remotely-sensed imagery as described below.
IV-A1 IEEE GRSS Data Fusion Contest 2022 (DFC2022) Dataset
The DFC2022 dataset [31] was released as part of the annual IEEE GRSS Data Fusion Contest. This dataset consists of images gathered in and around 19 urban areas from different regions in France. Each instance of this dataset contains a high-resolution RGB aerial image and its corresponding segmentation mask, both having approximately $\times 2000$ 2000 pixels and a spatial resolution of 50 cm per pixel, along with a Digital Elevation Model (DEM) with $\times 1000$ 1000 pixels at a spatial resolution of 100cm/pixel. For our experiments, we resample the DEM data to match the dimensions of the image and mask. The masks in this dataset contain 14 classes: “urban fabric”, “industrial, commercial, public, military, private and transport units”, “mine, dump and construction sites”, “artificial non-agricultural vegetated areas”, “arable land”, “permanent crops”, “pastures”, “complex and mixed cultivation patterns”, “orchards at the fringe of urban classes”, “forests”, “herbaceous vegetation associations”, “open spaces with little or no vegetation”, “wetlands”, and “water”. Examples of this dataset are shown in Figure 2.
Refer to captionRefer to captionRefer to caption
Figure 2: Example images (first row) of the DFC2022 dataset [31], their DEM data (second row), and the respective reference data (third row).
IV-A2 ISPRS Vaihingen and Potsdam Datasets
The Vaihingen and Potsdam datasets [32] were released for the 2D semantic labeling contest of the International Society for Photogrammetry and Remote Sensing (ISPRS). Both datasets consist of aerial imagery, Digital Surface Model (DSM) data, and label masks, as shown in Figure 3. The Vaihingen dataset contains 33 patches, with an average size of 2494×20642494$\times$ 20642494 × 2064 pixels. The aerial images have three bands (near-infrared, red, and green) with a spatial resolution of 9 cm per pixel. The Postdam dataset contains 38 tiles of $\times 6000$ 6000 pixels. The imagery consists of four bands (red, green, blue, and near-infrared) with a spatial resolution of 5 cm per pixel. The label masks in both datasets contain six classes: “impervious surfaces”, “building”, “low vegetation”, “tree”, “car”, and “clutter/background”.
Refer to captionRefer to captionRefer to caption
Figure 3: Example images (first row) of the Vaihingen and Potsdam datasets [32], their DSM data (second row), and the respective reference data (third row).
IV-B Implementation Details
We pre-process the aforementioned datasets by tiling them into non-overlapping $\times 256$ 256 patches that are used in all subsequent steps.
To evaluate a core-set of imagery, we train a U-Net [33] network with a ResNet-18 [28] backbone (that was pre-trained on the ImageNet dataset) on the core-set, and test this model on the held-out set separately for each experimental dataset. Importantly, our training and testing routine is fixed over the experiments, the only difference is in the subset (and size of subset) used to train the segmentation model.
All proposed methods [^1] The code is made publicly available at https://github.com/keillernogueira/data-centric-rs-classification/ are implemented using Pytorch. During training, we use the following hyper-parameters: 100 training epochs, AdamW [34] as optimizer, learning rate of 0.001, and batch size of 64 for the DFC2022 dataset and 32 for the other datasets.
For the LC/FD method, we let $m=770$ based on preliminary experiments with the DFC2022 dataset (we observed that FD outperformed LC for small subsets, while LC added value for larger datasets).
For the FA/CB method, we let $\lambda=0.5$ to equally weight the importance from the FA and FB methods.
IV-C Baselines
For all datasets, we compare the proposed techniques with two traditional baseline models: (i) Random, which selects the core set uniformly at random, and (ii) CoreSet [12, 21, 22], which selects samples to optimally cover the embedding space. Precisely, this approach iteratively expands the core-set by adding the data point that is farthest from its nearest neighbor in the current (core) set. In this case, we used the Euclidean distance in the activations of the last spatial pooling layer of the ResNet-18 [28], similar to the approach used in the FA and FD methods.
IV-D Experimental Protocol
For the DFC2022 dataset, 90% of the data originally released for the data fusion contest is made available to be ranked by the proposed core-set algorithms, while the remaining 10% is used for validation.
For the Vaihingen and Potsdam datasets, we follow the standard protocol commonly exploited in the literature [35]. Specifically, for the Vaihingen dataset, 11 images originally released for the contest are made available for the proposed core-set techniques, and 5 images (with IDs 11, 15, 28, 30, 34) are employed for validation. For the Potsdam dataset, 18 images released for the contest are made available for the proposed techniques, whereas 6 images (with IDs 02_12, 03_12, 04_12, 05_12, 06_12, 07_12) are used for validation.
For all datasets, the validation is only employed to assess the training of the U-Net models, after the selection of the core-set. The final evaluation of the trained U-Net models uses the original test set of each dataset. The overall performance of each method is measured by the mean Intersection over Union (mIoU) across all segmentation classes and averaged over three different model training runs.
V Experiments and Discussion
V-A Quantitative Results
[…]
Table I shows the results for each method across the three datasets and different core-set sizes. Overall, the proposed approaches outperform the baselines for all datasets, even when trained using 100% of the available data. Moreover, in most cases, the proposed approaches outperform both baselines even when using substantially fewer training examples. For instance, in the Vaihingen dataset, the FA/CB Hybrid at 10% achieved a higher mIoU score than both baselines on the top-ranked 25% of the data. This shows the ability of the proposed techniques to select representative examples (core-set) for training.
Furthermore, although the methods show varying degrees of effectiveness across different datasets, the image-based techniques tend to yield better performance with smaller training percentages (from 5 to 25%), producing good results within this range for the DFC2022 and Potsdam datasets. On the other hand, the label-based and combined approaches tend to produce good results for all training regimes, with the label-based approaches producing the best overall results for all datasets.
Finally, we also observe that each dataset has a different IoU convergence rate. DFC2022 has the most label noise (as evidenced by performance degradation of models in later splits), and the best-performing methods reached close to peak performance by utilizing only 10% of the data, noticeably outperforming training on 100% of the data. This rapid convergence suggests that for datasets with specific characteristics (e.g., redundancy, noise), a relatively small core-set can be as effective, or even more effective, than the full dataset. In contrast, on datasets with less label noise, such as Vaihingen and Potsdam, the performance continues to gradually improve as the number of training examples increases. However, even in this case, the introduced methods are able to outperform the baseline trained with 100% of the training data, demonstrating the importance of selecting a core set to deal with relevant issues such as noise, representativeness, and so on.
TABLE I: Results (mIoU) across different sized subsets of data (1%, 5%, 10%, 25%, 50%, 75%, and 100%) for DFC2022, Vaihingen, and Potsdam datasets. Underlined values indicate the results that outperformed the corresponding baselines per training percentage (statistically significant paired t-test at $\alpha=0.05$). Bold values represent the best results overall for the dataset.
Methods DFC 2022
Category 1% 5% 10% 25% 50% 75% 100%
Random - 10.02 $\pm$ 0.37 11.02 $\pm$ 0.23 11.72 $\pm$ 0.12 12.38 $\pm$ 0.29 12.68 $\pm$ 0.11 12.02 $\pm$ 0.11 12.29 $\pm$ 0.12
CoreSet [12, 21, 22] - 10.58 $\pm$ 1.54 10.58 $\pm$ 1.54 11.08 $\pm$ 0.40 11.08 $\pm$ 0.90 11.50 $\pm$ 0.52 12.17 $\pm$ 0.34 12.10 $\pm$ 0.85
Label Complexity (LC) Label-only 10.54 $\pm$ 0.63 12.04 $\pm$ 0.20 12.68 $\pm$ 0.20 12.65 $\pm$ 0.03 12.46 $\pm$ 0.16 11.73 $\pm$ 0.76 -
Feature Diversity (FD) Image-only 10.71 $\pm$ 0.71 11.57 $\pm$ 0.44 11.77 $\pm$ 0.49 11.96 $\pm$ 0.05 12.17 $\pm$ 0.29 11.86 $\pm$ 0.27 -
LC/FD Hybrid Both 10.27 $\pm$ 0.26 12.43 $\pm$ 0.25 12.09 $\pm$ 0.50 12.54 $\pm$ 0.14 12.24 $\pm$ 0.34 12.28 $\pm$ 0.35 -
Feature Activation (FA) Image-only 9.14 $\pm$ 0.54 10.85 $\pm$ 0.21 10.86 $\pm$ 0.37 11.63 $\pm$ 0.18 12.32 $\pm$ 0.12 12.45 $\pm$ 0.26 -
Class Balance (CB) Label-only 9.12 $\pm$ 0.52 10.69 $\pm$ 0.90 11.01 $\pm$ 0.58 10.58 $\pm$ 0.19 11.95 $\pm$ 0.22 11.99 $\pm$ 0.19 -
FA/CB Hybrid Both 10.13 $\pm$ 0.10 10.14 $\pm$ 0.25 10.82 $\pm$ 0.22 11.98 $\pm$ 0.61 11.62 $\pm$ 0.07 12.10 $\pm$ 0.22 -
Methods Vaihingen
Category 1% 5% 10% 25% 50% 75% 100%
Random - 36.20 $\pm$ 1.36 42.13 $\pm$ 0.52 50.72 $\pm$ 2.02 50.37 $\pm$ 2.97 55.82 $\pm$ 1.74 59.57 $\pm$ 0.64 59.95 $\pm$ 0.50
CoreSet [12, 21, 22] - 30.48 $\pm$ 1.49 32.65 $\pm$ 1.26 41.40 $\pm$ 0.54 45.68 $\pm$ 1.52 44.07 $\pm$ 3.85 49.17 $\pm$ 2.93 -
Label Complexity (LC) Label only 33.00 $\pm$ 4.03 45.80 $\pm$ 3.62 50.23 $\pm$ 4.33 53.75 $\pm$ 0.89 55.53 $\pm$ 0.41 58.70 $\pm$ 2.06 -
Feature Diversity (FD) Image only 33.67 $\pm$ 3.38 42.15 $\pm$ 0.57 44.87 $\pm$ 0.39 48.92 $\pm$ 3.60 56.07 $\pm$ 1.92 58.63 $\pm$ 0.50 -
LC/FD Hybrid Both 33.21 $\pm$ 3.50 40.45 $\pm$ 0.64 43.99 $\pm$ 0.42 51.15 $\pm$ 5.23 57.76 $\pm$ 0.46 58.97 $\pm$ 0.88 -
Feature Activation (FA) Image only 35.20 $\pm$ 1.22 43.61 $\pm$ 0.70 51.44 $\pm$ 0.69 53.44 $\pm$ 0.09 56.85 $\pm$ 0.60 60.58 $\pm$ 0.66 -
Class Balance (CB) Label only 40.92 $\pm$ 3.10 44.80 $\pm$ 1.23 51.03 $\pm$ 3.62 55.42 $\pm$ 0.83 58.36 $\pm$ 1.04 60.58 $\pm$ 0.30 -
FA/CB Hybrid Both 36.88 $\pm$ 1.39 49.06 $\pm$ 1.04 53.56 $\pm$ 0.25 56.92 $\pm$ 0.25 58.96 $\pm$ 0.67 60.43 $\pm$ 0.75 -
Methods Potsdam
Category 1% 5% 10% 25% 50% 75% 100%
Random - 60.89 $\pm$ 0.22 66.27 $\pm$ 0.88 68.21 $\pm$ 0.67 72.28 $\pm$ 0.66 74.56 $\pm$ 0.45 78.42 $\pm$ 0.33 78.00 $\pm$ 0.27
CoreSet [12, 21, 22] - 43.40 $\pm$ 7.30 42.69 $\pm$ 6.66 53.26 $\pm$ 6.27 68.85 $\pm$ 5.38 61.92 $\pm$ 1.28 64.98 $\pm$ 1.01 -
Label Complexity (LC) Label only 60.07 $\pm$ 2.33 68.68 $\pm$ 0.39 70.28 $\pm$ 0.45 75.12 $\pm$ 0.52 77.23 $\pm$ 0.79 79.15 $\pm$ 1.23 -
Feature Diversity (FD) Image only 59.78 $\pm$ 0.69 66.56 $\pm$ 0.35 69.37 $\pm$ 0.25 73.77 $\pm$ 0.54 75.42 $\pm$ 1.34 78.35 $\pm$ 1.21 -
LC/FD Hybrid Both 53.70 $\pm$ 6.05 68.25 $\pm$ 0.16 71.38 $\pm$ 0.34 74.87 $\pm$ 0.54 76.58 $\pm$ 0.56 78.72 $\pm$ 0.37 -
Feature Activation (FA) Image only 56.86 $\pm$ 1.47 65.75 $\pm$ 0.63 68.21 $\pm$ 0.58 72.89 $\pm$ 1.03 74.70 $\pm$ 0.52 79.58 $\pm$ 1.56 -
Class Balance (CB) Label only 49.89 $\pm$ 1.17 58.26 $\pm$ 0.39 61.94 $\pm$ 0.68 70.40 $\pm$ 0.30 75.29 $\pm$ 1.82 79.64 $\pm$ 0.48 -
FA/CB Hybrid Both 55.29 $\pm$ 1.49 61.72 $\pm$ 0.42 64.62 $\pm$ 0.40 71.06 $\pm$ 0.89 75.42 $\pm$ 0.78 78.49 $\pm$ 0.77 -
In addition to the gains in terms of performance, Table II reports the training time per epoch, in seconds, by a machine with an Intel Xeon E5-2695 v4 “Broadwell” CPU, 128GB of RAM memory, Nvidia P100 ”Pascal” GPU with 16GB of memory, under an 11.2 CUDA version, and Red Hat Enterprise Linux system release 8.2. As all models are trained for 100 epochs, this allows us to contextualize computational savings from the core-set methods. For example, with the DFC2022 dataset, the best-performing result (using 10% of the data) trained for 2 days and 2 hours less than the 100% training baseline while achieving a better test set performance. It is important to highlight that the core-set selection processing time is of the order of magnitude of one epoch with 100% of the data or less, making it negligible compared to the full 100-epoch training. Additionally, faster training time allows for more epochs and/or more extensive hyperparameter tuning, further enhancing model optimization.
TABLE II: Training time per epoch (in seconds)
Dataset 1% 5% 10% 25% 50% 75% 100%
DFC2022 40.65 107.34 194.62 428.14 810.97 1191.38 1981.11
Vaihingen 8.17 15.42 16.73 21.39 39.81 57.34 76.22
Potsdam 13.05 20.2 28.58 54.74 97.51 141.29 185.2
V-B Qualitative Results
To facilitate the analysis and comparison of the proposed methods’ outputs, we include visualizations of the average generated rankings, as can be seen in Figure 4. To generate these visualizations, the rankings produced by the different introduced methods are first averaged by patch position and then sorted, making the highest-ranking patches across the methods appear at the top. Additionally, the standard deviation is calculated to capture the variability of the assigned ranks and provide insight into the approaches’ consistency. In addition to these visualizations, to allow for a better analysis, we also report the Kendall Tau correlations between each pair of methods in Figure 5 and provide examples of the most and least frequently selected instances across all introduced methods in Figure 6.
(a) DFC2022
(b) Vaihingen
(c) Potsdam
Figure 4: Visualizations of the proposed methods’ rankings. The line represents the average rank position for each patch across all proposed approaches, while the shaded area represents the standard deviation.
(a) DFC2022
(b) Vaihingen
(c) Potsdam
Figure 5: Correlation of methods according to Kendall Tau coefficient. A high correlation value means that the methods produce similar rankings.
For all datasets, it is possible to observe that the approaches exhibit notable consistency, frequently assigning more importance to a specific (core) set of high-complexity examples that are consistently selected across the proposed methods. Similarly, such approaches also tend to agree on the least important examples, assigning lower scores to low-complexity patches, indicating that the explored datasets contain a subset of non-representative or noisy instances that either contribute minimally to the overall performance or, in some cases, may even degrade it. Furthermore, this level of agreement between the proposed techniques can be further observed in the correlation plots, wherein several methods show substantial correlation (particularly for the DFC2022 and Potsdam datasets), suggesting that they can capture underlying dataset patterns (such as the core sets). Overall, the ability to select the core set, along with the identification of less valuable examples, highlights the robustness and efficiency of the proposed techniques in distinguishing between high- and low-quality data, thereby resulting in better performance and training time. This is also qualitatively demonstrated in Figure 6: examples that are consistently highly ranked by the proposed methods show clear imagery with artifacts that are of a certain visual and semantic complexity (as illustrated by the corresponding label maps). A few of these images allow learning the appearance and spatial relation of several classes at once. On the other hand, images that are consistently rejected show very homogeneous scenes (such as large water bodies or parking lots) with neither much visual variation nor complex semantic content.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(a) DFC2022
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(b) Vaihingen
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(c) Potsdam
Figure 6: Examples of the highest-ranked (first three columns) and lowest-ranked instances (last three columns) considering the average ranking of all proposed methods.
VI Conclusions
In this paper, we introduce and benchmark six basic core-set selection approaches for remote sensing image segmentation based on distinct premises - which rely on imagery only, labels only, or a combination of each - thereby establishing a general and comprehensive baseline for future works. The proposed methods are able to consistently and effectively select the most important subset of examples (i.e., core-set) while filtering out non-representative and/or noisy instances.
Experiments are conducted using three high-resolution remote sensing datasets with very distinct properties: (i) IEEE GRSS Data Fusion Contest 2022 (DFC2022) dataset [31], consisting of very high-resolution visible spectrum images and Digital Elevation Model imagery, and (ii) Vaihingen and Potsdam datasets [32], both composed of high-resolution multispectral images and normalized Digital Surface Model data.
Experimental results demonstrate the effectiveness and computational efficiency of the proposed techniques, which consistently outperform the baselines. Notably, on the DFC2022 dataset, the proposed approaches outperform the baseline trained on 100% of the data while using only 10% of the available examples. Similarly, on the Vaihingen and Potsdam datasets, the same superior performance is achieved using just 75% of the data. Overall, the core-set selection not only enhances the performance of the deep learning models but also substantially reduces training time.
In summary, this work addresses a crucial gap in the literature and demonstrates the potential of core-set selection in advancing remote sensing image segmentation as well as data creation and labeling. The presented conclusions open opportunities towards: (i) the integration of core-set selection with other advanced techniques, such as self-supervised learning and foundation models, and (ii) a more efficient and effective exploitation of both existing and new datasets for a better understanding of the Earth’s surface, an essential characteristic for most applications.